OpenStudy (anonymous):

does this form a ratio or proportion?8/12,12/18

OpenStudy (anonymous):

yes, they do. Both are a rario of 2 to 3,

OpenStudy (anonymous):

does it form a ratio or proportion

OpenStudy (anonymous):

8/12 = 12/18

OpenStudy (anonymous):

Ratio, proportion, they mean the same.

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

Fractions is another term that applies. They can be reduced to lowest terms.

OpenStudy (anonymous):

both reduced to lower terms = 2/3

OpenStudy (anonymous):

ohh ok i kinda get it now

OpenStudy (anonymous):

as a ratio they 1:1

OpenStudy (anonymous):

ohhh ok thx ur pretty ku

OpenStudy (anonymous):

8/12 = 12/18 = 2/3.

OpenStudy (anonymous):

a ratio is a comparison of two numbers with the same units while a proportion is two equal ratios. So each is a ratio and together they make a correct proportion